# Displacive Jahn–Teller transition in NaNiO2

### v2

@@ 1 Introduction
Figure 1: Crystal structures of (a) $C2/m$ monoclinic and (c) $R\bar{3}m$ rhombohedral NaNiO2. (b) and (d) show the Ni-Ni distances in these two structures. For the monoclinic structure, short Ni-Ni distances are in grey and long Ni-Ni distances are in green. For the rhombohedral structure, all nearest-neighbour distances are equal and shown in grey.
@@ 2 Results
Selected results of the Rietveld analysis are shown in Figure 2. The volume per formula unit is larger in the rhombohedral than the monoclinic structure [Figure 2(b)]. There is positive thermal expansion in both the monoclinic and rhombohedral phases. In the monoclinic phase, NiO6 octahedral volume is essentially temperature-independent [inset of Figure 2(b)] and the volume increase is entirely driven by NaO6 octahedral expansion. Through the phase transition, there is a discontinuous increase in unit cell volume per formula unit, primarily due to a $\sim$1.2jump in NiO6 octahedral volume (compared with a $\sim$0.6increase for NaO6 octahedra) from the fully monoclinic to fully rhombohedral phases. In the rhombohedral phase both the NiO6 and NaO6 octahedra expand on heating at around the same rate.
@@
Figure 3: (a) Experimental neutron PDFs for NaNiO2 at 293 K, 450 K, and 500 K, in the $r$-range where the nearest-neighbour Ni-O peaks occur. Peaks are labelled for the 293 K dataset in blue text and 500 K dataset in green text. (b,c) Example fits to the PDF and Bragg diffraction data at 500 K using the supercell fitting (experimental data: blue; fitted data: orange; difference pattern; green).
@@
Figure 5: (a) Ni-O bond distributions at 293 K (top), 450 K (middle), and 500 K (bottom), as obtained via big box analysis of neutron PDF data, for the smallest to largest Ni-O bond in each octahedron; vertical lines are the bond lengths from Rietveld refinement of the neutron data reported in Ref. 27. (b) Histogram of bond length distortion index 43 for each NiO6 octahedron at 293 K, 450 K, and 500 K. (c) Probability function of distortion [Equation 1] at 293 K, 450 K, and 500 K. (b) and (c) are obtained using VanVleckCalculator 6, 44.
Other evidence for the absence of Jahn–Teller distortion in the local structure of rhombohedral NaNiO2 can be obtained by considering two metrics for the octahedral distortion. The bond length distortion index (BLDI) 43, Figure 5(b), quantifies the deviation of bond lengths from the average for a given polyhedron and is often used to quantify Jahn–Teller distortions 45, 46, 27. For a big box model where thermal effects are modelled by deviation of atoms from their average positions (rather than using atomic displacement parameters, as is conventional for Rietveld refinement), we would expect the BLDI to increase with heating in the absence of any electronic or magnetostructural transitions. The average BLDI for all NiO6 octahedra is 0.0547 and 0.0575 (to 3 significant figures) at 293 K and 450 K, respectively, in the monoclinic phase, before decreasing to 0.0458 at 500 K. This is clearly shown by the decreasing distribution of bond lengths in Figure 5(a). In contrast, the average BLDI continuously increases with heating for the NaO6 octahedra, increasing from 0.0260 at 293 K, to 0.0352 at 450 K, and finally 0.0450 at 500 K (see Figure S10). This trend in the BLDI for NiO6 octahedra can be explained by the disappearance of a local Jahn–Teller distortion at $T_{\mathrm{JT}}$. We can also quantify the distortion using the parameter $\rho_{0}$, where $\rho_{0}^{2}=Q_{2}^{2}+Q_{3}^{2}$. Figure 5(c) shows $P(\rho_{0})$, the probability of octahedra having a given value of $\rho_{0}$, defined as:
@@
where $f(\rho_{0})$ is a histogram of $\rho_{0}$ for all NiO6 octahedra. This parameterisation shows maximum probability at $\rho_{0}\approx 0$ at 500 K, compared with large non-zero magnitudes at 293 K and 450 K, which also indicates the absence of a Jahn–Teller distortion in the rhombohedral phase.
@@
XAS at the Ni K edge was performed, with analysis focusing on the extended X-ray absorption fine structure (EXAFS) in real-space. This data, $\chi(r)$, is analogous to a partial PDF of Ni-O bond lengths, but with the significant difference that there is a phase shift resulting in peaks in $\chi(r)$ being down-shifted by around 0.5 Å. It is therefore a useful supplement to the PDF data. The room-temperature $\chi(r)$ data presented here resembles the EXAFS on NaNiO2 reported previously 47.
@@
Models were obtained using FEFF 48 and refined against the $\chi(r)$ data over different ranges. Two different R(Å) ranges were considered; one with just the Ni-O shell R(Å) = 0.5 – 2.0 Å, and the other R(Å) = 0.5 – 3.1 Å to also include the Ni-Ni interatomic distances. Models with both 1 or 2 Ni-O and Ni-Ni distances were considered and we describe these using curly braces enclosing the comma-separated values for the number of lengths for each bond; for instance a model with a single Ni-O distance in the reduced range is described as {1,0}, whereas a model with two Ni-O bond lengths and two Ni-Ni bond lengths is described as {2,2}. Five models were used, over both ranges: {1,0}, {2,0}, {1,1}, {2,1}, and {2,2}. Models in which the first number is 1 correspond to JT-undistorted, and models where the first element is 2 are compatible with a JT distortion.
@@ 3 Discussion
Our studies of the local Ni3+ environment, using neutron PDF and Ni K edge EXAFS, suggest the absence of local static Jahn–Teller-distortions for $T>T_{\mathrm{JT}}$. Further, while a dynamic Jahn–Teller distortion 4 has previously been suggested for LiNiO2 50, 51, 52, by analogy with similar proposals for LaMnO3 53, such a dynamic Jahn–Teller distortion would also be resolvable via our neutron PDF and EXAFS measurements. These observations, taken together with the increase in unit cell volume for NaNiO2, suggest that the JT transition in NaNiO2 is best characterised as displacive rather than order-disorder 13. This is different to the observations of an order-disorder transition in LaMnO3 18, 15, 14, 19, and the proposed order-disorder transition for NaNiO2 13. To our knowledge, only in LiNiO2 17 has prior evidence for a displacive Jahn–Teller transition been put forward, using AIMD and inferences from changes in strain.
@@
As in previous reports we observe a first order phase transition in NaNiO2 with the monoclinic and rhombohedral phases coexisting over a finite temperature range. The observed discontinuous increase in unit cell volume from the monoclinic to the rhombohedral structures [Figure 2] is also consistent with a first-order phase transition. The positive thermal expansion through the transition is consistent, via the Clausius-Clapeyron equation, with our previous finding 27 of $dT_{\mathrm{JT}}/dP>0$. We note that in the displacive transition in SrTiO3 54 a positive $dT_{\mathrm{a}}/dP$ is also observed. In contrast, a negative thermal expansion is found at the transition temperature in LaMnO3 55 and ferroelectric PbTiO3 56, 57, both order-disorder transitions, with $dT_{c}/dP<0$ 58, 59. The displacive transition may therefore be related to the sign of $dT_{\mathrm{JT}}/dP$.
@@
In NaNiO2, the high-temperature phase will be enthalpically unfavourable and hence the phase transition will be entropically-driven, according to the free energy $\Delta G=\Delta H-T\Delta S$, where $\Delta H$ is the enthalpy change, $T$ is temperature, and $\Delta S$ is the entropy change. There are two relevant contributions to the entropy: configurational and vibrational. We find the configurational entropy of orbital disorder for LiNiO2 and NaNiO2 to be subextensive (see Appendix A), meaning it does not increase with increasing system size 60. A subextensive configurational entropy has also been found for an Ising model on an elastic triangular lattice 61. Consequently the entropy term is dominated by dynamic vibrational effects which disfavour local Jahn–Teller distortions. This is likely a contributing factor in the displacive JT transition we have observed, and is in contrast to high-temperature LaMnO3 where, by analytic and Monte-Carlo calculations using the Potts model, there is extensive configurational entropy of orbital disorder 20, 21 and hence an order-disorder transition.
@@
An alternative hypothesis we considered for the disappearance of Jahn–Teller distortions in high-temperature NaNiO2 was the possibility for delocalisation of the electrons resulting in a change in the energy landscape of the system. This is by analogy to the insulator-to-metal transition (IMT) previously observed in the perovskite nickelates 62, 63. However, while variable-temperature conductivity, $\sigma(T)$, data from Ref. 64 [Figure S6], shows an increase in conductivity by several orders of magnitude on heating, there is no change of sign of $d\sigma(T)/dT$. This is similar to the conductivity data for LaMnO3 53 which does not show a displacive transition, and so we did not consider this hypothesis further.
@@
There are a number of other differences between NaNiO2 and LaMnO3 which could affect the Jahn–Teller transition. For instance, the edge-sharing octahedral connectivity in NaNiO2 results in much larger intra-octahedral angular distortion compared with the corner-sharing octahedra in LaMnO3. This difference in connectivity is a consequence of the triangular Ni network in NaNiO2. Given the constraint that axes of elongation of adjacent JT-distorted Ni atoms cannot point at the same O anion, there are no possible long-range JT-disordered configurations 30, which means that this triangular Ni network cannot host a Potts-type Jahn–Teller disorder such as that proposed for LaMnO3 20, 21, 16. Another difference is that $d^{4}$ Mn3+ is generally found to be JT-distorted, whereas nominally $d^{7}$ Ni3+ is susceptible to other types of electronic configuration such as charge disproportionation 38, 39, 36, 37, 40; this may be due to the smaller electronegativity difference between Ni-O, compared with Mn-O, which has been argued to cause charge-transfer insulating 65 behaviour in many nickelates 66, 63, 67. Attempting to deconvolute the role of these different factors will require further study on the transition in other Jahn–Teller-distorted materials, along with theoretical and computational work.
@@ Methods
Variable-temperature, ambient-pressure synchrotron X-ray diffraction was performed using the I11 instrument 68, 69 at Diamond Light Source ($\lambda=0.824110$) using the Mythen-2 position-sensitive detector, with a data collection time of $\sim$10 seconds. The sample was contained in a 0.5 mm diameter glass capillary, sealed inside a glovebox with epoxy (Loctite Double Bubble). The sample was heated at a rate of 12 K/min from 322 K to 796 K, with periodic measurements.
Data were analysed by sequential Rietveld refinement 41 using the software package Topas 7 70. A pseudo-Voigt peak function was used, and background was fitted using a Chebyshev polynomial with 20 terms. A small correction for preferred orientation was applied, using the March-Dollase model 71, 72.
@@
Samples for variable-temperature nuclear magnetic resonance (VT-NMR) were packed into 4 mm ZrO2 magic angle spinning (MAS) rotors in an Ar filled glovebox, and fitted with ZrO2 caps. 23Na chemical shifts were calibrated using solid NaCl as an external secondary reference (7.21 ppm relative to 1 M NaCl(aq) at 0.0 ppm). 23Na MAS NMR spectra were acquired using a Bruker Avance IIIHD spectrometer (v0[1H] = 500.13 MHz, v0[23Na] = 132.46 MHz, v0 [207Pb] = 104.26 MHz) spectrometer, with a Bruker 4 mm HX probe and 14 kHz MAS. Projection magic-angle turning and a phase-adjusted sideband separation (pj-MATPASS) pulse sequence was used to obtain the isotropic 23Na spectra 73.
Static VT-NMR samples were measured in the same spectrometer using a Bruker HX static probe. A Hahn-echo pulse sequence with $\pi/4$ = 2.05 $\mu$s optimized on solid NaNiO2 was used. Temperature was calibrated through a NMR shift thermometer compound Pb(NO3)2, based on the known temperature-dependence of the isotropic chemical shift of Pb(NO3)2 74. 207Pb NMR spectra were acquired with the same probe at the same temperature values, using a Hahn-echo pulse sequence. The isotropic chemical shift values for Pb(NO3)2 were obtained by fitting the data using the SOLA fitting program within Topspin 4.1.4.
@@
Variable-temperature total-scattering neutron diffraction was performed on the NOMAD instrument 75 at the Spallation Neutron Source, Oak Ridge National Laboratory, USA. NaNiO2 was stored under Ar in a sealed NMR tube (2 cm sample height, 5 mm outer diameter) for the measurements. Heating was performed using a furnace, at a rate of 5K per minute. The sample was measured on heating at 293 K, 450 K, 500 K, and after cooling at 316 K. The neutron diffraction data and the refined crystal structure was previously published in Ref. 27. In the Fourier transform to real space, a sliding $Q_{\mathrm{max}}$ as a function of $r$ is used, with $Q_{\mathrm{max}}=40$-1 at low-$r$ and gradually decreasing $Q_{\mathrm{max}}$ beyond that, with $Q_{\mathrm{max}}=20$-1 above $r=25$.
@@
Small-box pair distribution function analysis, also known as real-space Rietveld refinement, was performed using Topas 6 70. Data were fitted in the range 1.0 Å to 10 Å in real space only. The starting structures were the normal $C2/m$ and $R\bar{3}m$ structures used commonly in the literature for the low- and high-temperature phases of NaNiO2.
Pair distribution function analysis was performed via a “big box” approach used Topas 7 70, following a modified version of the method introduced in Ref. 76. A supercell model was refined against both the Bragg data from NOMAD bank 4 (as defined in Ref. 75) and the experimental neutron pair distribution function. In this refinement, atomic coordinates were refined until convergence following the Rietveld algorithm 41. During initial iterations, restraints were applied to prevent large atomic movements, but these were removed in later stages of the refinement. Upon convergence, each atom was randomly shifted by a distance less than or equal to 0.1% of the unit cell size ($\sim$0.045 Å) in each spatial dimension to help minimisation. This typically occurred several hundred times. The final configuration was the converged structure with the best fit quality. Our approach differs from the more commonly-used approach in RMCProfile 77. Data were fitted in the range 0.5 Å to 20 Å in real space.
A $16\times 9\times 3$ supercell of a pseudo-orthorhombic cell (space group: $P1$; $a\approx 2.846$; $b\approx 5.321$; $c\approx 15.703$; $\beta=\gamma=90^{\circ}$; $\alpha=90^{\circ}$ for the 500 K supercell and $\sim$88 for the $T<T_{\mathrm{JT}}$ supercells) was used for the refinements. This NaNiO2 unit cell was obtained via transformation from the rhombohedral cell using Equation S5 in Supplementary Information. The supercell has edge lengths $\sim$45 Å and contains 10368 atoms. In this analysis, the thermal parameters for all atoms were fixed at low values ($B_{\mathrm{iso}}=0.01$2) such that thermal effects are modelled by the deviation of each site from its ideal position. To ensure that chemically reasonable models were obtained, soft restraints were included. The relative weighting of restraints against data was set to ensure refinements were not dominated by restraints. The SI includes refinements performed in the absence of restraints to check their influence on the conclusions (Section S3.9). The main soft restraint was the calculated bond valence sum 78 (as an analog for oxidation state), $V_{\mathrm{calc}}$, of cations deviates from its expected value (Na: +1; Ni: +3), an established approach 79 in PDF analysis 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 76, 99, 100, 101, 102. Here, $V_{\mathrm{calc}}$ is calculated as follows:
@@
The supercells obtained from refinement were then analysed using VanVleckCalculator 6, 44, a Python 3 103 code which is based on PyMatGen 104.

### v4

@@ 1 Introduction
Figure 1: Crystal structures of (a) $C2/m$ monoclinic and (c) $R\bar{3}m$ rhombohedral NaNiO2. (b) and (d) show the Ni-Ni distances (grey for non-elongated; green for elongated) and JT-elongated Ni-O distances (pink) in these two structures. In the rhombohedral structure there are no elongated Ni-Ni or Ni-O bonds.
@@ 2 Results
Selected results of the Rietveld analysis are shown in Figure 2, with further information in SI Section 2. The volume per formula unit is larger in the rhombohedral than the monoclinic structure [Figure 2(b)]. There is positive thermal expansion in both the monoclinic and rhombohedral phases. In the monoclinic phase, NiO6 octahedral volume is essentially temperature-independent [inset of Figure 2(b)] and the volume increase is entirely driven by NaO6 octahedral expansion. Through the phase transition, there is a discontinuous increase in unit cell volume per formula unit, primarily due to a $\sim$1.2jump in NiO6 octahedral volume (compared with a $\sim$0.6increase for NaO6 octahedra) from the fully monoclinic to fully rhombohedral phases. In the rhombohedral phase both the NiO6 and NaO6 octahedra expand on heating at around the same rate.
@@
Figure 3: (a) Experimental neutron PDFs for NaNiO2 at 293 K, 450 K, and 500 K, in the $r$-range where the nearest-neighbour Ni-O peaks occur. Peaks are labelled for the 293 K dataset in blue text and 500 K dataset in green text. The 500 K data is vertically offset from the low-temperature data for ease of distinguishing the changes compared with the lower-$T$ datasets. Grey dashed horizontal lines occur at $D(r)=0$ for the lower two temperatures (lower) and 500 K (upper). (b,c) Example fits to the PDF and Bragg diffraction data at 500 K using the supercell fitting (experimental data: blue; fitted data: orange; difference pattern; green).
@@
To this end, a rhombohedral NaNiO2 unit cell with centred orthorhombic setting was prepared, consisting of 6 formula units and obtained via transformation from the normal rhombohedral cell using Equation S5. A $16\times 9\times 3$ supercell of this unit cell was used as the starting point for big box refinements, with lattice dimensions $\sim$45 Å and containing 10368 atoms; see Methods for further details. The atomic positions in this model were then refined against the experimental data. Here, empirical bond valence sum 43 restraints (discussed in Methods) were applied to the atomic positions to ensure physical behaviour, although qualitatively equivalent results are obtained without using these restraints [SI Section S3.9]. The big box refinements were run several times to ensure repeatability [SI Section S3.11].
@@
Figure 5: (a) Ni-O bond distributions at 293 K (top), 450 K (middle), and 500 K (bottom), as obtained via big box analysis of neutron PDF data, for the smallest to largest Ni-O bond in each octahedron; vertical lines are the bond lengths from Rietveld refinement of the neutron data reported in Ref. 27. (b) Histogram of bond length distortion index 44 for each NiO6 octahedron at 293 K, 450 K, and 500 K. (c) Probability function of distortion [Equation 1] at 293 K, 450 K, and 500 K. (b) and (c) are obtained using VanVleckCalculator 6, 45.
Other evidence for the absence of Jahn–Teller distortion in the local structure of rhombohedral NaNiO2 can be obtained by considering two metrics for the octahedral distortion. The bond length distortion index (BLDI) 44, Figure 5(b), quantifies the deviation of bond lengths from the average for a given polyhedron and is often used to quantify Jahn–Teller distortions 46, 47, 27. For a big box model where thermal effects are modelled by deviation of atoms from their average positions (rather than using atomic displacement parameters, as is conventional for Rietveld refinement), we would expect the BLDI to increase with heating in the absence of any electronic or magnetostructural transitions. The average BLDI for all NiO6 octahedra is 0.0547 and 0.0575 (to 3 significant figures) at 293 K and 450 K, respectively, in the monoclinic phase, before decreasing to 0.0458 at 500 K. This is clearly shown by the decreasing distribution of bond lengths in Figure 5(a). In contrast, the average BLDI continuously increases with heating for the NaO6 octahedra, increasing from 0.0260 at 293 K, to 0.0352 at 450 K, and finally 0.0450 at 500 K (see Figure S10). This trend in the BLDI for NiO6 octahedra can be explained by the disappearance of a local Jahn–Teller distortion at $T_{\mathrm{JT}}$. We can also quantify the distortion using the parameter $\rho_{0}$, where $\rho_{0}^{2}=Q_{2}^{2}+Q_{3}^{2}$. Figure 5(c) shows $P(\rho_{0})$, the probability of octahedra having a given value of $\rho_{0}$, defined as:
@@
where $f(\rho_{0})$ is a histogram of $\rho_{0}$ for all NiO6 octahedra. This parameterisation shows maximum probability at $\rho_{0}\approx 0$ at 500 K, compared with large non-zero magnitudes at 293 K and 450 K, which also indicates the absence of a local Jahn–Teller distortion in the rhombohedral phase. This result that the most probable $\rho_{0}$ substantially decreases at 500 K, compared with lower temperatures, is resilient to other ways of running the big box refinement, for instance with different starting configurations [SI Section 3.7], changes in the BVS restraints [SI Section S3.9], and multiple runs with the same starting parameters [SI Section 3.11].
@@
XAS at the Ni K edge was performed, with analysis focusing on the extended X-ray absorption fine structure (EXAFS) in real-space. This data, $\chi(r)$, is analogous to a partial PDF of Ni-O bond lengths, but with the significant difference that there is a phase shift resulting in peaks in $\chi(r)$ being down-shifted by around 0.5 Å. It is therefore a useful supplement to the PDF data. The room-temperature $\chi(r)$ data presented here resembles the EXAFS on NaNiO2 reported previously 48.
@@
Models were obtained using FEFF 49 and refined against the $\chi(r)$ data over different ranges. Two different R(Å) ranges were considered; one with just the Ni-O shell R(Å) = 0.5 – 2.0 Å, and the other R(Å) = 0.5 – 3.1 Å to also include the Ni-Ni interatomic distances. Models with both 1 or 2 Ni-O and Ni-Ni distances were considered and we describe these using curly braces enclosing the comma-separated values for the number of lengths for each bond; for instance a model with a single Ni-O distance in the reduced range is described as {1,0}, whereas a model with two Ni-O bond lengths and two Ni-Ni bond lengths is described as {2,2}. Five models were used, over both ranges: {1,0}, {2,0}, {1,1}, {2,1}, and {2,2}. Models in which the first number is 1 correspond to JT-undistorted, and models where the first element is 2 are compatible with a JT distortion.
@@ 3 Discussion
Our studies of the local Ni3+ environment, using neutron PDF and Ni K edge EXAFS, suggest the absence of local static Jahn–Teller-distortions for $T>T_{\mathrm{JT}}$. Further, while a dynamic Jahn–Teller distortion 4 has previously been suggested for LiNiO2 51, 52, 53, by analogy with similar proposals for LaMnO3 54, such a dynamic Jahn–Teller distortion would also be resolvable via our neutron PDF and EXAFS measurements. These observations suggest that the JT transition in NaNiO2 is better characterised as displacive rather than order-disorder 13. This is different to the observations of an order-disorder transition in LaMnO3 18, 15, 14, 19, and the proposed order-disorder transition for NaNiO2 13. To our knowledge, only in LiNiO2 17 has prior evidence for a displacive Jahn–Teller transition been put forward, using AIMD and inferences from changes in strain.
@@
As in previous reports we observe a first order phase transition in NaNiO2 with the monoclinic and rhombohedral phases coexisting over a finite temperature range. The observed discontinuous increase in unit cell volume from the monoclinic to the rhombohedral structures [Figure 2] is also consistent with a first-order phase transition. The positive thermal expansion through the transition is consistent, via the Clausius-Clapeyron equation, with our previous finding 27 of $dT_{\mathrm{JT}}/dP>0$. We note that in the displacive transition in SrTiO3 55 a positive $dT_{\mathrm{a}}/dP$ is also observed. In contrast, a negative thermal expansion is found at the transition temperature in LaMnO3 56 and ferroelectric PbTiO3 57, 58, both order-disorder transitions, with $dT_{c}/dP<0$ 59, 60. The displacive transition may therefore be related to the sign of $dT_{\mathrm{JT}}/dP$.
@@
In NaNiO2, the high-temperature phase will be enthalpically unfavourable and hence the phase transition will be entropically-driven, according to the free energy $\Delta G=\Delta H-T\Delta S$, where $\Delta H$ is the enthalpy change, $T$ is temperature, and $\Delta S$ is the entropy change. There are two relevant contributions to the entropy: configurational and vibrational. We find the configurational entropy of orbital disorder for LiNiO2 and NaNiO2 to be subextensive (see Appendix A), meaning it does not increase with increasing system size 61. A subextensive configurational entropy has also been found for an Ising model on an elastic triangular lattice 62. Consequently the entropy term is dominated by dynamic vibrational effects which disfavour local Jahn–Teller distortions. This is likely a contributing factor in the displacive JT transition we have observed, and is in contrast to high-temperature LaMnO3 where, by analytic and Monte-Carlo calculations using the Potts model, there is extensive configurational entropy of orbital disorder 20, 21 and hence an order-disorder transition.
@@
An alternative hypothesis we considered for the disappearance of Jahn–Teller distortions in high-temperature NaNiO2 was the possibility for delocalisation of the electrons resulting in a change in the energy landscape of the system. This is by analogy to the insulator-to-metal transition (IMT) previously observed in the perovskite nickelates 63, 64. However, while variable-temperature conductivity, $\sigma(T)$, data from Ref. 65 [Figure S6], shows an increase in conductivity by several orders of magnitude on heating, there is no change of sign of $d\sigma(T)/dT$. This is similar to the conductivity data for LaMnO3 54, which would suggest the electronic conductivity is not closely related to the order-disorder/displacive nature of the transition.
@@
There are a number of other differences between NaNiO2 and LaMnO3 which could affect the Jahn–Teller transition. For instance, the edge-sharing octahedral connectivity in NaNiO2 results in much larger intra-octahedral angular distortion compared with the corner-sharing octahedra in LaMnO3. This difference in connectivity is a consequence of the triangular Ni network in NaNiO2. Given the constraint that axes of elongation of adjacent JT-distorted Ni atoms cannot point at the same O anion, there are no possible long-range JT-disordered configurations 30, which means that this triangular Ni network cannot host a Potts-type Jahn–Teller disorder such as that proposed for LaMnO3 20, 21, 16. Another difference is that $d^{4}$ Mn3+ is generally found to be JT-distorted, whereas nominally $d^{7}$ Ni3+ is susceptible to other types of electronic configuration such as charge disproportionation 38, 39, 36, 37, 40; this may be due to the smaller electronegativity difference between Ni-O, compared with Mn-O, which has been argued to cause charge-transfer insulating 66 behaviour in many nickelates 67, 64, 68. Attempting to deconvolute the role of these different factors will require further study on the transition in other Jahn–Teller-distorted materials, along with theoretical and computational work.
Whether or not a JT transition is displacive or order-disorder may be important in some battery cathode materials. This will be particularly apparent if the JT transition occurs below room-temperature, as it does in, for instance, the NaNixCo1-xO2 solid solution for $x\leq 0.8$ 65 and the spinel LiMn2O4 69. In these cases, the presence or absence of local JT distortions will result in different ionic mobility, and will impact the crystallographic changes that occur with electrochemical cycling. The implications of this on electrochemistry should be studied in future works.
@@ Methods
Variable-temperature, ambient-pressure synchrotron X-ray diffraction was performed using the I11 instrument 70, 71 at Diamond Light Source ($\lambda=0.824110$) using the Mythen-2 position-sensitive detector, with a data collection time of $\sim$10 seconds. The sample was contained in a 0.5 mm diameter glass capillary, sealed inside a glovebox with epoxy (Loctite Double Bubble). The sample was heated at a rate of 12 K/min from 322 K to 796 K, with periodic measurements.
Data were analysed by sequential Rietveld refinement 41 using the software package Topas 7 72. A pseudo-Voigt peak function was used, and background was fitted using a Chebyshev polynomial with 20 terms. A small correction for preferred orientation was applied, using the March-Dollase model 73, 74.
@@
Samples for variable-temperature nuclear magnetic resonance (VT-NMR) were packed into 4 mm ZrO2 magic angle spinning (MAS) rotors in an Ar filled glovebox, and fitted with ZrO2 caps. 23Na chemical shifts were calibrated using solid NaCl as an external secondary reference (7.21 ppm relative to 1 M NaCl(aq) at 0.0 ppm). 23Na MAS NMR spectra were acquired using a Bruker Avance IIIHD spectrometer (v0[1H] = 500.13 MHz, v0[23Na] = 132.46 MHz, v0 [207Pb] = 104.26 MHz) spectrometer, with a Bruker 4 mm HX probe and 14 kHz MAS. Projection magic-angle turning and a phase-adjusted sideband separation (pj-MATPASS) pulse sequence was used to obtain the isotropic 23Na spectra 75.
Static VT-NMR samples were measured in the same spectrometer using a Bruker HX static probe. A Hahn-echo pulse sequence with $\pi/4$ = 2.05 $\mu$s optimized on solid NaNiO2 was used. Temperature was calibrated through a NMR shift thermometer compound Pb(NO3)2, based on the known temperature-dependence of the isotropic chemical shift of Pb(NO3)2 76. 207Pb NMR spectra were acquired with the same probe at the same temperature values, using a Hahn-echo pulse sequence. The isotropic chemical shift values for Pb(NO3)2 were obtained by fitting the data using the SOLA fitting program within Topspin 4.1.4.
@@
Variable-temperature total-scattering neutron diffraction was performed on the NOMAD instrument 77 at the Spallation Neutron Source, Oak Ridge National Laboratory, USA. NaNiO2 was stored under Ar in a sealed NMR tube (2 cm sample height, 5 mm outer diameter) for the measurements. Heating was performed using a furnace, at a rate of 5K per minute. The sample was measured on heating at 293 K, 450 K, 500 K, and after cooling at 316 K. The neutron diffraction data and the refined crystal structure was previously published in Ref. 27. In the Fourier transform to real space, a sliding $Q_{\mathrm{max}}$ as a function of $r$ is used, with $Q_{\mathrm{max}}=40$-1 at low-$r$ and gradually decreasing $Q_{\mathrm{max}}$ beyond that, with $Q_{\mathrm{max}}=20$-1 above $r=25$.
@@
Small-box pair distribution function analysis, also known as real-space Rietveld refinement, was performed using Topas 6 72. Data were fitted in the range 1.0 Å to 10 Å in real space only. The starting structures were the normal $C2/m$ and $R\bar{3}m$ structures used commonly in the literature for the low- and high-temperature phases of NaNiO2.
Pair distribution function analysis was performed via a “big box” approach used Topas 7 72, following a modified version of the method introduced in Ref. 78. A supercell model was refined against both the Bragg data from NOMAD bank 4 (as defined in Ref. 77) and the experimental neutron pair distribution function. In this refinement, atomic coordinates were refined until convergence following the Rietveld algorithm 41. During initial iterations, restraints were applied to prevent large atomic movements, but these were removed in later stages of the refinement. Upon convergence, each atom was randomly shifted by a distance less than or equal to 0.1% of the unit cell size ($\sim$0.045 Å) in each spatial dimension to help minimisation. This typically occurred several hundred times. The final configuration was the converged structure with the best fit quality. Our approach differs from the more commonly-used approach in RMCProfile 79. Data were fitted in the range 0.5 Å to 20 Å in real space.
A $16\times 9\times 3$ supercell of a pseudo-orthorhombic cell (space group: $P1$; $a\approx 2.846$; $b\approx 5.321$; $c\approx 15.703$; $\beta=\gamma=90^{\circ}$; $\alpha=90^{\circ}$ for the 500 K supercell and $\sim$88 for the $T<T_{\mathrm{JT}}$ supercells) was used for the refinements. This NaNiO2 unit cell was obtained via transformation from the rhombohedral cell using Equation S5 in Supplementary Information. The supercell has edge lengths $\sim$45 Å and contains 10368 atoms. In this analysis, the thermal parameters for all atoms were fixed at low values ($B_{\mathrm{iso}}=0.01$2) such that thermal effects are modelled by the deviation of each site from its ideal position. To ensure that chemically reasonable models were obtained, soft restraints were included. The relative weighting of restraints against data was set to ensure refinements were not dominated by restraints. The SI includes refinements performed in the absence of restraints to check their influence on the conclusions (Section S3.9). The main soft restraint was the calculated bond valence sum 43 (as an analog for oxidation state), $V_{\mathrm{calc}}$, of cations deviates from its expected value (Na: +1; Ni: +3), an established approach 80 in PDF analysis 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 78, 100, 101, 102, 103. Here, $V_{\mathrm{calc}}$ is calculated as follows:
@@
The supercells obtained from refinement were then analysed using VanVleckCalculator 6, 45, a Python 3 104 code which is based on PyMatGen 105.